# “Stupid robot, I want to speak to a human!” User Frustration Detection in Task-Oriented Dialog Systems

### v1

@@ Appendix B Dialog Breakdown Detection Features
Feature name Computation methodology
Semantic paraphrase of user’s req. MA of the cosine similarity between pairs of utterances $u_{t-1}$ and $u_{t}$
Semantic repetition of system’s resp. MA of the cosine similarity between pairs of utterances $s_{t-1}$ and $s_{t}$
Semantic coherence of user’s req. and system’s resp. MA of the cosine similarity between utterances $s_{t-1}$ and $u_{t}$
Syntactic paraphrase of user’s req. MA of the Jaccard index between sets of words of utterances $u_{t-1}$ and $u_{t}$
Syntactic repetition of system’s resp. MA of the Jaccard index between sets of words of utterances $s_{t-1}$ and $s_{t}$
Syntactic coherence of user’s req. and system’s resp. MA of the Jaccard index between sets of words of utterances $s_{t-1}$ and $u_{t}$
Length of user’s requests MA of the utterance length $u_{t}$
Length of system’s response MA of the utterance length $s_{t}$
Length of full conversation Number of characters in observed dialog $H$
Number of turns in conversation Number of pairs $(u_{t},s_{t})$

### v2

@@ 4 Results
Macro-$F_{1}$
Sentiment Analysis Hartmann et al. (2023)
RoBERTa-Sent-FC 0.34
RoBERTa-Sent-LU 0.65
Emotion Detection
DistilBERT-EmoWoZ-FC Huang (2024) 0.56
DistilBERT-EmoWoZ-LU 0.60
DistilRoBERTa-Emo-FC Hartmann (2022) 0.40
DistilRoBERTa-Emo-LU 0.44
Dialog Breakdown Detection
DBD+LogReg Bodigutla et al. (2020) 0.71
Rule-Based Approach
Keyword Matching 0.41
LLM-based ICL Approach
GPT-4o-zero-shot-LU 0.67
GPT-4o-two-shot-LU 0.75
Llama-3.1-405B-zero-shot-LU 0.63
Llama-3.1-405B-two-shot-LU 0.75
GPT-4o-zero-shot-FC 0.86
GPT-4o-two-shot-FC 0.84
Llama-3.1-405B-zero-shot-FC 0.83
Llama-3.1-405B-two-shot-FC 0.87
Table 2: Comparison of UF detections performed on last user utterance (LU) and full conversations (FC).
@@ Appendix B Dialog Breakdown Detection Features
Feature name Computation methodology Semantic paraphrase of user’s req. MA of the cosine similarity between pairs of utterances $u_{t-1}$ and $u_{t}$ Semantic repetition of system’s resp. MA of the cosine similarity between pairs of utterances $s_{t-1}$ and $s_{t}$ Semantic coherence of user’s req. and system’s resp. MA of the cosine similarity between utterances $s_{t-1}$ and $u_{t}$ Syntactic paraphrase of user’s req. MA of the Jaccard index between sets of words of utterances $u_{t-1}$ and $u_{t}$ Syntactic repetition of system’s resp. MA of the Jaccard index between sets of words of utterances $s_{t-1}$ and $s_{t}$ Syntactic coherence of user’s req. and system’s resp. MA of the Jaccard index between sets of words of utterances $s_{t-1}$ and $u_{t}$ Length of user’s requests MA of the utterance length $u_{t}$ Length of system’s response MA of the utterance length $s_{t}$ Length of full conversation Number of characters in observed dialog $H$ Number of turns in conversation Number of pairs $(u_{t},s_{t})$